t^2 + 2(5.0 m/s) /−9.80 m/s^2


1
Expert's answer
2020-12-03T10:23:56-0500

I think that given things are not enough to compute but i will simplify it:

"t^2" + "\frac{2\times(5.0\frac{m}{s})}{-9.80\frac{m}{s^2}}"="t^2" - 1.02s. --> this is the thing that we reached maximum simplify.